A minimum height is to be prescribed for eligibility to
government services such that 75.80% of the young men will have a
fair chance of coming up to that standard. The height of young
men is normally distributed with mean 62 inches and standard
deviation 3.5 inches. Determine the minimum specification.
1
Expert's answer
2020-04-20T18:43:55-0400

"Given \; that, μ=62, σ=3.5,\\
Let\; Z_{1}=\frac{X-62}{3.5}, then\\
 P(Z>Z_{1})=0.758\\
0.5+P(-Z_{1}<Z<0)=0.758\\
P(-Z_{1}<Z<0)=0.258\\
-Z_{1}=0.7
\implies Z_{1}=-0.7\\
\therefore -0.7=\frac{X-62}{3.5}\\
X=(-0.7\times 3.5)+62=59.55\\
\text{the minimum specification is}\;59.55\; inches"